# On the combinatorics of tableaux — Classification of lattices underlying Schensted correspondences

### v1

@@ 6. Global characterization
If $T_{xy}$ and $T_{x+1,y+1}$ are both assigned, then either $T_{x+1,y}$ or $T_{x,y+1}$ or both are assigned and $w(T_{xy})=w(T_{x+1,y+1})$. As a consequence, if there is a sequence of assigned locations $T_{x+i,y+i}$ for $0\leq i\leq n$ then all of these locations have the same value of $w$. (This set of locations has the same value of $h$ and $v$ values $x+y+2i$.)
The left and right sides of $J$
We start by analyzing the right side of $J$, the elements assigned to $T_{xy}$ where $y\geq x$.
Definition 6.15.
If there is a maximum value of $y$ for which $T_{0y}$ is assigned, we define that value to be $y_{\textup{max}}$. In that case, $y_{\textup{max}}\geq 1$. If $T_{0y}$ is assigned for all $y\geq 0$, we define $y_{\textup{max}}=\infty$.
Note that we have not proven that $y_{\textup{max}}$ is the largest value of $y$ for which some $T_{xy}$ is defined. We use the convention that $i<\infty$ for any (finite) integer $i$.
@@
$w(T_{0y})=m+y\delta^{+}\textup{\qquad\qquad for all }1\leq y\leq y_{\textup{max}}$ (31)
$y_{\textup{max}}$ is finite iff $\delta^{+}<0$. If $y_{\textup{max}}$ is finite, then
$y_{\textup{max}}=-m/\delta^{+}-1.$ (32)
which is equivalent to $w(T_{0,y_{\textup{max}}})=-\delta^{+}$ and
$m=-(y_{\textup{max}}+1)\delta^{+}.$ (33)
In particular, $\delta^{+}$ must divide $m$
@@
Because $T_{10}$ is assigned, unique-cover-modularity implies that $T_{1y}$ is assigned for all $y\leq y_{\textup{max}}$. This implies that for any $1\leq y\leq y_{\textup{max}}$, none of the $T_{1y}$ are upward orphans of the the corresponding $T_{0y}$, as $T_{1,y-1}$ is assigned. Thus the constraints on $w$ due to applying lem. 5.26 to $T_{0y}$ for $1\leq y<y_{\textup{max}}$ are $2w(T_{0y})=w(T_{0,y-1})+w(T_{0,y+1})$, which collectively imply (31).
If $y_{\textup{max}}$ is finite, then by definition $T_{0,y_{\textup{max}}+1}$ is not assigned and the constraint on $w$ for $T_{0,y_{\textup{max}}}$ is $2w(T_{0,y_{\textup{max}}})=w(T_{0,y_{\textup{max}}-1})$. Applying (31) and lem. 6.14 gives $2(m+y_{\textup{max}}\delta^{+})=m+(y_{\textup{max}}-1)\delta^{+}$, which implies (32).
Conversely, assume $\delta^{+}\geq 0$. Then (32) cannot be satisfied by any finite $y_{\textup{max}}\geq 1$. Hence by the preceding paragraph, there can be no finite $y_{\textup{max}}$. ∎
@@
For all $x<y$, applying (31) and lem. 6.14 gives
$w(T_{xy})=w(T_{0,y-x})=m+(y-x)\delta^{+}=m+h\delta^{+}$ (34)
where h=yxh=y-x as defined above.
Definition 6.18.
Define $c_{\textup{max}}$ as the maximum integer such that for all $0\leq i\leq c_{\textup{max}}$, $T_{ii}$ is assigned, or if all $T_{ii}$ are assigned, then $c_{\textup{max}}=\infty$.
Note that we have not proved that there is no $T_{jj}$ assigned with $j>c_{\textup{max}}$. The mnemonic for $c_{\textup{max}}$” is that it is the index of the top element of the centerline extending upward from $T_{00}$.
@@
By unique-cover-modularity, all $T_{xy}$ are assigned for all $x,y$ with $x<y$, $x\leq c_{\textup{max}}$, and $y\leq y_{\textup{max}}$.
@@
If $y\geq x+2$, $T_{x+1,y-1}$ is not assigned, and $T_{x,y-1}$, $T_{xy}$, and $T_{x-1,y}$ are assigned, then $T_{x+1,y}$ is not assigned. (See fig. 7(a) for a diagram of the configuration.)
(a) lem. 6.20
(b) lem. 6.21
Figure 7. Configurations for lem. 6.20 and 6.21, showing the weights of the elements relative to $\alpha=w(T_{xy})$. Solid dots indicate assigned elements, empty dots indicate unassigned elements, and crosses can be either.
@@
The constraint for $T_{xy}$ is
$2w(T_{xy})=w(T_{x,y-1})+w(T_{x-1,y})+w(T_{x+1,y})+w(T_{x,y+1})$
where the third term on the right is present iff $T_{x+1,y}$ is an upward orphan of $T_{xy}$ and the fourth term on the right is present iff $T_{x,y+1}$ is an upward orphan of $T_{xy}$. If we define $\alpha=w(T_{xy})$, then this simplifies to
$2\alpha=(\alpha-\delta^{+})+(\alpha-\delta^{+})+w(T_{x+1,y})+w(T_{x,y+1}).$
Since both $w(T_{x+1,y})$ and $w(T_{x,y+1})$ must be positive, both $T_{x+1,y}$ and $T_{x,y+1}$ must be either unassigned or not upward orphans of $T_{xy}$. Since by hypothesis $T_{x+1,y-1}$ is unassigned, $T_{x+1,y}$ would be an upward orphan if it was assigned. Thus $T_{x+1,y}$ is not assigned. ∎
@@
If $y\geq x+1$, $T_{x-1,y+1}$ is not assigned, and $T_{x-1,y}$, $T_{xy}$, and $T_{x,y-1}$ are assigned, then $T_{x,y+1}$ is not assigned. (See fig. 7(b) for a diagram of the configuration.)
@@
This is proved by a left-right reversal of the proof of lem. 6.20. ∎
Lemma 6.22.
Given the values of $y_{\textup{max}}$ and $c_{\textup{max}}$, the set of assigned $T_{xy}$ of $J^{+}$ with $x\leq y$ is as shown in the corresponding diagram in fig. 8.
(a) $c_{\textup{max}}=\infty$ and $y_{\textup{max}}=\infty$
(b) $c_{\textup{max}}=\infty$ and $y_{\textup{max}}$ finite
(c) $c_{\textup{max}}$ finite and $y_{\textup{max}}=\infty$
(d) $c_{\textup{max}}$ finite, $y_{\textup{max}}$ finite, and $y_{\textup{max}}>c_{\textup{max}}$
(e) $c_{\textup{max}}$ finite, $y_{\textup{max}}$ finite, and $y_{\textup{max}}=c_{\textup{max}}$
(f) $c_{\textup{max}}$ finite, $y_{\textup{max}}$ finite, and $y_{\textup{max}}<c_{\textup{max}}$
Figure 8. The possible sets of assignments on the left side of $J^{+}$. Solid dots, heavy lines, and shading indicate assigned $T_{\bullet\bullet}$’s. Empty dots and light lines indicate unassigned $T_{\bullet\bullet}$’s.
Proof.
In each diagram of fig. 8, the $T_{\bullet\bullet}$ on the centerline that are shows as assigned and unassigned is determined by the corresponding hypothesis regarding $c_{\textup{max}}$. Similarly, the $T_{\bullet\bullet}$ on the $y$ axis that are shows as assigned and unassigned is determined by the corresponding hypothesis regarding $y_{\textup{max}}$.
The assigned $T_{\bullet\bullet}$ represented by the shaded areas are determined by lem 6.19.
The series of unassigned $T_{\bullet\bullet}$ extending upward and to the right (increasing $y$) from $T_{c_{\textup{max}}+1,c_{\textup{max}}+1}$ are determined by lem. 6.20. The series of unassigned $T_{\bullet\bullet}$ extending upward and to the left (increasing $x$) from $T_{0,y_{\textup{max}}+1}$ are determined by lem. 6.21.
In case (D), the non-allocation $T_{c_{\textup{max}}+1,y_{\textup{max}}+1}$ is determined by lem. 6.14, since neither $T_{c_{\textup{max}}+1,y_{\textup{max}}}$ $T_{c_{\textup{max}},y_{\textup{max}}+1}$ is assigned. ∎
The analysis of the left side of $J$, the elements assigned to $T_{xy}$ where $x\geq y$, is the left-right reversal to the analysis of the right side.
Definition 6.23.
If there is a maximum value of $x$ for which $T_{x0}$ is assigned, we define that value to be $x_{\textup{max}}$. In that case, $x_{\textup{max}}\geq 1$. If $T_{x0}$ is assigned for all $x\geq 0$, we define $x_{\textup{max}}=\infty$.
Lemma 6.24.
Given the values of $x_{\textup{max}}$ and $c_{\textup{max}}$, the set of assigned $T_{xy}$ of $J^{+}$ with $x\geq y$ is the left-right reversal of diagram in fig. 8 for $y_{\textup{max}}=-x_{\textup{max}}$ and $c_{\textup{max}}$.
The global structure of $J$
Thus, the overall structure of the allocation of the lower ranks of $J^{+}$ can be described by two letters, the first being the case of fig. 8 that describes the lower part of the left-hand side (where $h\leq 0$) and the second letter being the case that describes the lower part of the right-hand side (where $h\geq 0$). Since swapping the two letters corresponds to left-right reversing the allocations, leaving the poset $J$ isomorphic, without loss of generality, we list only letter pairs where the first is alphabetically before or the same as the second.
Most combinations of left and right cases determine the entire assignment of $J^{+}$ as the allocations and non-allocations in the two cases determine the allocation of the entirety of $J^{+}$. E.g. (AA) specifies that all $T_{\bullet\bullet}$ are assigned. But also (CE) specifies that all $T_{\bullet,c_{\textup{max}}+1}$ are unassigned, which requires that no $T_{\bullet y}$ is assigned for $y>c_{\textup{max}}$. On the other hand, (AB) does not directly specify the allocation of $T_{xy}$ when $y_{\textup{min}}\leq x<y$.
Lemma 6.25.
The possible structures of the allocation of the lower ranks of $J^{+}$ are (AA), (AB), and (CE).
Proof.
Of necessity, the left side and the right side have the same value of $c_{\textup{max}}$, and in particular it is either finite or infinite. Thus the possible structures are composed of letters A and B (if $c_{\textup{max}}=\infty$), or of letters C, D, E, and F (if $c_{\textup{max}}$ is finite).
The following cases lead to contradictions.
  1. ($B$):
    The allocation of $T_{y_{\textup{max}},y_{\textup{max}}}$ and $T_{y_{\textup{max}}+1,y_{\textup{max}}+1}$ violates lem. 6.14, since neither $T_{y_{\textup{max}}+1,y_{\textup{max}}}$ nor $T_{y_{\textup{max}},y_{\textup{max}}+1}$ is assigned.
  2. ($C$),
    (CD), and (DD): The allocation of $T_{c_{\textup{max}},c_{\textup{max}}}$, $T_{c_{\textup{max}}+1,c_{\textup{max}}}$, and $T_{c_{\textup{max}},c_{\textup{max}}+1}$ but not $T_{c_{\textup{max}}+1,c_{\textup{max}}+1}$ violates unique-cover-modularity.
  3. (CF):
    The constraint for $T_{y_{\textup{max}}+1,y_{\textup{max}}}$ (which is assigned on the left side) is
    $2w(T_{y_{\textup{max}}+1,y_{\textup{max}}})=w(T_{y_{\textup{max}}+1,y_{\textup{max}}-1})+w(T_{y_{\textup{max}},y_{\textup{max}}})+w(T_{y_{\textup{max}}+1,y_{\textup{max}}+1})$
    because $T_{y_{\textup{max}}+1,y_{\textup{max}}+1}$ is assigned and $T_{y_{\textup{max}},y_{\textup{max}}+1}$ is not, making $T_{y_{\textup{max}}+1,y_{\textup{max}}+1}$ an upward orphan of $T_{y_{\textup{max}}+1,y_{\textup{max}}}$. That equation can be simplified to
    $2(m+\delta^{-})=(m+2\delta^{+})+m+w(T_{y_{\textup{max}}+1,y_{\textup{max}}+1})$
    which is impossible since $w(T_{y_{\textup{max}}+1,y_{\textup{max}}+1})$ is positive.
  4. (DE),
    (DF), (EE), (EF), and (FF): The constraint for $T_{11}$ (which must be assigned) is
    $2w(T_{11})=w(T_{10})+w(T_{01})$
    That simplifies to $2m=(m+\delta^{-})+(m+\delta^{+})$, so $\delta^{-}+\delta^{+}=0$. But in this case both $x_{\textup{max}}$ and $y_{\textup{max}}$ are finite, so both $\delta^{+}$ and $\delta^{-}$ must be negative, which is a contradiction.
Thus only (AA), (AB), and (CE) are possibilities. ∎
Lemma 6.26.
If the structure of the allocation $J^{+}$ is (AA), then the structure of $J$ is:
  1. (1)
    All $T_{\bullet\bullet}$ are assigned.
  2. (2)
    $N=1$ and $m=r$.
  3. (3)
    $w(T_{xy})=r$ for all $x,y$.
Thus,
Case 6.27.
(para. 2(2)) $J$ is isomorphic to the quadrant of the plane, $\{(x,y)\,|\,\textup{$x$ and $y$ are integers $\geq 0$}\}$, and $L$ is $\mathbb{Y}$, Young’s lattice, the lattice of partitions, with weighting $r$ times the conventional weighting of 1.
Proof.
By the hypothesis (AA), all $T_{\bullet\bullet}$ are assigned.
The constraint for $T_{11}$ is
$2w(T_{11})=w(T_{10})+w(T_{01})$
That simplifies to $2m=(m+\delta^{-})+(m+\delta^{+})$, so $\delta^{-}+\delta^{+}=0$. In this case $x_{\textup{max}}=y_{\textup{max}}=\infty$, so both $\delta^{+}$ and $\delta^{-}$ must be nonnegative, and thus $\delta^{-}=\delta^{+}=0$.
By (30), $N=1$. Thus $T_{00}=B_{1}$.
By (29), $m=w(B_{1})=r$. Since $\delta^{-}=\delta^{+}=0$, by (31) and its left-right reverse, $w(T_{\bullet\bullet})=r$. ∎
Lemma 6.28.
If the structure of the allocation $J^{+}$ is (AB), then the structure of $J$ is:
  1. (1)
    $T_{xy}$ is assigned iff $y\leq x+1$.
  2. (2)
    $N=2$.
  3. (3)
    $w(T_{xy})=2r$ if $y\leq x$.
  4. (4)
    $w(T_{xy})=r$ if $y=x$.
Thus,
Case 6.29.
[…]
Proof.
By the hypothesis (AB), all $T_{xy}$ with $x\geq y$ are assigned, all $T_{xy}$ with $x\leq y\leq y_{\textup{max}}$ are assigned, and all $T_{x,y_{\textup{max}}+1}$ with $x\leq y_{\textup{max}}-1$ are unassigned.
If $y_{\textup{max}}\geq 2$, the constraint for $T_{11}$ (which must be assigned) is
$2w(T_{11})=w(T_{10})+w(T_{01})$
because neither $T_{21}$ or $T_{12}$ are upward orphans of $T_{11}$. That simplifies to $2m=(m+\delta^{-})+(m+\delta^{+})$, so $\delta^{-}+\delta^{+}=0$. In this case $x_{\textup{max}}=\infty$ and $y_{\textup{max}}$ is finite, so $\delta^{+}$ must be negative and $\delta^{-}$ must be nonnegative, which is a contradiction. Thus $y_{\textup{max}}=1$, which implies $m=-2\delta^{+}$.
The allocations and non-allocations that are known so far are as shown in fig. 9(a).
(a) Lem. 6.28, first step.
(b) Lem. 6.28, second step.
Figure 9. Solid dots indicate assigned elements, empty dots indicate unassigned elements, and crosses can be either.
If $T_{02}$ is not assigned, then the constraint for $T_{11}$ is
$2w(T_{11})=w(T_{10})+w(T_{01})$
That simplifies to $2m=(m+\delta^{-})+(m+\delta^{+})$, so $\delta^{-}+\delta^{+}=0$. But by the same argument as above, this is a contradiction. Thus $T_{12}$ is assigned.
The constraint for $T_{11}$ is
$2w(T_{11})=w(T_{10})+w(T_{01})+w(T_{12})$
That simplifies to $2m=(m+\delta^{-})+(m+\delta^{+})+(m+\delta^{+})$ and since $m=-2\delta^{+}$, $\delta^{-}=0$.
The allocations and non-allocations that are known so far are as shown in fig. 9(b).
This analysis can be iterated, examining the constraints for $T_{ii}$ for increasing $i\geq 1$, giving the complete allocation of $J^{+}$ shown in fig. 10.
Figure 10. Solid dots indicate assigned elements and empty dots indicate unassigned elements.
The constraint for $T_{00}$ is
$2w(T_{00})=w(B_{N})=w(B_{N-1})+w(T_{10})+w(T_{01})$
which simplifies to $2Nr=(N-1)r+(m+\delta^{-})+(m+\delta^{+})$. Since $m=Nr=-2\delta^{+}$ and $\delta^{-}=0$, this reduces to $r=-\delta^{+}$, and thus $N=2$ and $m=2r$. ∎
Lemma 6.30.
If the structure of the allocation $J^{+}$ is (CE), then the structure of $J$ is:
  1. (1)
    $r=\alpha k$ for some positive integers $\alpha,k$.
  2. (2)
    $T_{xy}$ is assigned iff $y<k$.
  3. (3)
    $N=1$ and $m=r$.
  4. (4)
    $w(T_{xy})=\alpha(k+x-y)$ for all $0\leq x$ and $0\leq y<k$.
Thus,
Case 6.31.
(para. 2(4)) $J$ is isomorphic to the $k$-row strip, $\{(x,y)\,|\,\textup{$x$ and $y$ are integers, $0\leq x$ and $0\leq y<k$}\}$, and $L$ is $\mathbb{Y}_{k}$, the $k$-row Young’s lattice, the lattice of partitions with at most $k$ parts, with weighting $\alpha>0$ times the conventional weighting $w(T_{xy})=k+x-y$.
Proof.
By the hypothesis (CE), $c_{\textup{max}}=y_{\textup{max}}$. All $T_{xy}$ are assigned for $0\leq y<y_{\textup{max}}$. Also, all $T_{x,y_{\textup{max}}+1}$ are unassigned which constrains all $T_{xy}$ with $y\geq y_{\textup{max}}$ to be unassigned.
The constraint for $T_{11}$ (which must be assigned) is
$2w(T_{11})=w(T_{10})+w(T_{01})$
That simplifies to $2w=(w+\delta^{-})+(w+\delta^{+})$, so $\delta^{-}+\delta^{+}=0$. In this case $x_{\textup{max}}=\infty$ and $y_{\textup{max}}$ is finite, $\delta^{+}$ is a negative integer and $\delta^{-}=-\delta^{+}$ is a positive integer. By (30), $N=1$. Thus $T_{00}=B_{1}$. By (29), $m=w(B_{1})=r$.
Define $k=y_{\textup{max}}+1$ and $\alpha=-\delta^{+}=\delta^{-}$. Thus by (33), $r=m=k\alpha$. By (31) and its left-right reverse, $w(T_{xy})=\alpha(k+x-y)$. ∎

### v2

@@ 6. Global characterization
Not considering the order of the terms, simple enumeration shows the decompositions of $1$ as a sum of reciprocals as in (33) are:
$\frac{1}{1}$ $+\frac{1}{n}$ $+(-\frac{1}{n})\textup{\quad for any integer $n\geq 1$}$
$\frac{1}{1}$ $+\frac{1}{\infty}$ $+\frac{1}{\infty}$
$\frac{1}{2}$ $+\frac{1}{2}$ $+\frac{1}{\infty}$
$\frac{1}{2}$ $+\frac{1}{3}$ $+\frac{1}{6}$
$\frac{1}{2}$ $+\frac{1}{4}$ $+\frac{1}{4}$
$\frac{1}{3}$ $+\frac{1}{3}$ $+\frac{1}{3}$
The possible cases for $P$
@@
Since
  1. (1)
    the possible decompositions of 1 as a sum of reciprocals are listed in lem. 6.18,
  2. (2)
    NN must be positive,
  3. (3)
    given the left-right symmetry of $P$, we conventionally choose ${x_{L}}\geq{y_{L}}$, and
  4. (4)
    $T_{10}$ and $T_{01}$ are allocated, so $\Delta^{-}\neq 1$ and $\Delta^{+}\neq 1$,
the possible cases for $P$ are limited to:
$\Delta^{-}$ $x_{L}$} $\Delta^{+}$ $y_{L}$} NN Lattice
n-n for $\geq 2 \infty nn$ $n$ 1 $\mathbb{Y}_{n}$
$\infty$ $\infty$ $\infty$ $\infty$ 1 $\mathbb{Y}$
$\infty$ $\infty$ 2 2 2 $\mathbb{SY}$
3 3 2 2 6 none (lem. 6.22
6 6 2 2 3 none (lem. 6.24
6 6 3 3 2 none (lem. 6.21)
4 4 2 2 4 none (lem. 6.23)
4 4 4 4 2 none (lem. 6.21)
3 3 3 3 3 none (lem. 6.21)
Remark 6.20.
If we do not require $\hat{0}$, then we remove the restriction $N>0$ in lem. 6.19. That allows the additional case $\Delta^{-}=\Delta^{+}=2$, and $N=\infty$. Developing that case leads to the Fomin lattices para. 2(11).
Lemma 6.21.
It is not possible that $x_{L}$} and $y_{L}$} are not $\infty$ and are >2>2.
@@
If $x_{L}$} and $y_{L}$} are not $\infty$ and are >2>2, then $T_{20}$ and $T_{02}$ are allocated and neither $T_{21}$ nor $T_{12}$ are upward orphans of $T_{11}$. Thus the constraints of lem. 5.28 for $T_{11}$ simplifies to $0=\delta^{-}+\delta^{+}$. But since $x_{L}$} and $y_{L}$} are finite, $\delta^{-}<0$ and $\delta^{+}<0$, which is a contradiction. ∎
@@
It is not possible that ${x_{L}}=3$ and ${y_{L}}=2$.
Proof.
Figure 14. The case ${x_{L}}=3$ and ${y_{L}}=2$: Elements are labeled with their weights in multiples of $m$. Coverings marked with a stroke are upward orphan coverings. Double vertical strokes connect weights that are equal by lem. 5.27(2). Empty circles are $T_{\bullet\bullet}$ that are not allocated. The constraint on $T_{44}$ must be violated.
Applying the constraints of lem. 5.28 to $T_{\bullet\bullet}$ in increasing order of $v$ straightforwardly determines the allocations, non-allocations, and weights shown in fig. 14. But of necessity, $w_{44}=w_{33}=m$, which violates the constraint on $T_{44}$. ∎
@@
It is not possible that ${x_{L}}=4$ and ${y_{L}}=2$.
Proof.
Figure 15. The case ${x_{L}}=4$ and ${y_{L}}=2$: Elements are labeled with their weights in multiples of $m$. Coverings marked with a stroke are upward orphan coverings. Double vertical strokes connect weights that are equal by lem. 5.27(2). Empty circles are $T_{\bullet\bullet}$ that are not allocated. The constraint on $T_{22}$ must be violated.
Applying the constraints of lem. 5.28 to $T_{\bullet\bullet}$ in increasing order of $v$ straightforwardly determines the allocations, non-allocations, and weights shown in fig. 15. But of necessity, $w_{22}=w_{11}=m$, which violates the constraint on $T_{22}$. ∎
@@
It is not possible that ${x_{L}}=6$ and ${y_{L}}=2$.
@@
Figure 16. The case ${x_{L}}=6$ and ${y_{L}}=2$: Elements are labeled with their weights in multiples of $m$. Coverings marked with a stroke are upward orphan coverings. Double vertical strokes connect weights that are equal by lem. 5.27(2). Empty circles are $T_{\bullet\bullet}$ that are not allocated. The constraint on $T_{22}$ must be violated.
Applying the constraints of lem. 5.28 to $T_{\bullet\bullet}$ in increasing order of $v$ straightforwardly determines the allocations, non-allocations, and weights shown in fig. 16. But of necessity, $w_{22}=w_{11}=m$, which violates the constraint on $T_{22}$. ∎
@@
For the first three cases of lem. 6.19, $L$ must be:
  1. (1)
    for $\Delta^{-}=-n$ and $\Delta^{+}=n$, $L$ is isomorphic to $\mathbb{Y}_{n}$ with weights that are $m/n$ times the canonical weights (para. 2(4) and fig. 3),
  2. (2)
    for $\Delta^{-}=\Delta^{+}=\infty$, $L$ is isomorphic to $\mathbb{Y}$ with weights that are $m$ times the canonical weights (para. 2(2) and fig. 1),
  3. (3)
    for $\Delta^{-}=\infty$ and $\Delta^{+}=2$, $L$ is isomorphic to $\mathbb{SY}$ with weights that are $m/2$ times the canonical weights (para. 2(3) and fig. 2),
with $P$ being the poset of join-irreducibles of $L$ with the corresponding weights.
@@
In all of these cases, the values $w_{B_{i}}$, $w_{x0}$, and $w_{0y}$ are determined (relative to $m$) by $x_{L}$}, $y_{L}$}, and $N$. Which other elements of $T_{\bullet\bullet}$ are allocated and their weights are straightforwardly determined by the constraints of lem. 5.28 on them (working in order of increasing $v$), and necessarily result in the known lattices listed in the table. ∎